# Polaritonic Control of Blackbody Infrared Radiative Dissociation

### v2

@@ Abstract
Vibrational strong light-matter coupling offers a promising approach for controlling chemical reactivity with infrared microcavities. While recent research has examined potential mechanisms for this phenomenon, many important questions remain, including what type of reactions can be modified and to what extent this modification can be achieved. In this study, we explore the dynamics of Blackbody Infrared Radiative Dissociation (BIRD) in microcavities under weak and strong light-matter interaction regimes. Using a Master equation approach, we simulate the effects of infrared field confinement and vibrational strong coupling on BIRD rates for diatomic molecules. We present a framework explaining how infrared microcavities influence BIRD kinetics, highlighting the importance of overtone transitions in the process. Our findings outline the conditions under which significant enhancement or mild suppression of BIRD rates can be achieved, offering insights into the practical limitations and new strategies for controlling chemistry within infrared resonators.
@@ I   Introduction
Chemical reaction control is critical in applications ranging from industrial catalysis to medicine. Conventional methods for manipulating reaction rates involve varying temperature, pressure, or catalyst composition. An unconventional approach for controlling thermal reaction kinetics has been recently reported [1, 2, 3, 4] employing infrared (IR) microcavities [5] consisting of two moderate-quality mirrors separated by a distance $L_{C}$ of the order of IR wavelengths resonant with typical molecular vibrations [6]. When a collection of molecules with sufficiently large IR oscillator strength is embedded in a resonant microcavity, vibrational strong coupling occurs as signaled by the emergence of hybrid quasiparticles (vibrational polaritons) consisting of a superposition of molecular and electromagnetic modes of the confined device [7, 8, 9]. Recent theoretical research has probed potential mechanisms for the observed polariton effects on chemical reaction rates[10, 11, 12, 13, 14, 15, 16, 4, 17, 18, 19]. Still, several fundamental questions remain [20, 21], including what types of reactions can be controlled with an optical microcavity and to what extent this control can be exerted. In this study, we investigate gas-phase unimolecular dissociation activated by the absorption of microcavity blackbody infrared radiation [22] under weak and strong coupling with a suitable material.
@@
Blackbody infrared radiative dissociation (BIRD) was proposed by Perrin in 1913 as a mechanism for gas phase unimolecular dissociation [23]. However, experimental observations in the 1920s ruled BIRD out in favor of the collisional activation mechanism [24, 25, 26]. Several decades later, the BIRD hypothesis was proven appropriate at sufficiently low pressures[27, 28, 29]. Since then, BIRD has become a valuable tool for investigating the thermal dissociation kinetics of molecular ions and clusters and has been observed in several simple and complex systems, including ion clusters, transition-metal complexes, and biopolymers [22].
@@
The thermal radiation spectrum heavily influences BIRD rates. Given that free space and microcavity electromagnetic density of states and thermal radiation density can be significantly different [5], microcavity BIRD is expected to proceed with different rates compared to free space BIRD. However, the extent of this change, whether suppression or enhancement can be achieved, and if strong light-matter coupling introduces new possibilities relative to the weak coupling regime remain open questions. Our study provides answers to these questions. Our quantitative analysis focuses on the dissociation of diatomic molecules induced by multiphoton or multipolariton absorption in microcavities, but several qualitative conclusions are generalizable to polyatomic systems.
@@
The primary aim of this study is to characterize thermal radiative dissociation in microcavities. We employ a Master equation treatment to investigate this process in (i) free space, (ii) under weak coupling with a nearly empty microcavity, and (iii) under weak coupling with a polaritonic material. Our results give upper bounds for BIRD enhancement and suppression in weakly and strongly coupled microcavities and identify overtone transitions as crucial ingredients for the observed modulation of BIRD. In Sec. II, we describe our methodology, Sec. III provides the main results and related discussion, and Sec. IV summarizes our main results and conclusions.
@@ II   Methods
Figure 1: (a) Illustration of the Morse potential describing the vibrational states and transitions of the HF molecule. (b) and (c): A scheme illustrating an HF molecule inside a planar cavity and its dispersion under the weak coupling regime, respectively. (d) and (e): A scheme illustrating an HF molecule inside a planar cavity with the host material strongly coupled to the cavity and the dispersion of polariton states, respectively.
Master Equation for diatomic BIRD
Our analysis of BIRD in microcavities employs similar tools to those successfully applied to investigate this process in free space [28, 30, 31, 22]. These include a Markovian Master equation describing the time-evolution of vibrational quantum state populations with energy below the dissociation threshold defined below. Rotational degrees of freedom are ignored since they are weakly coupled to vibrational transitions, and the rotational dynamics are essentially classical at the temperatures of interest to BIRD. We also neglect Doppler and lifetime broadening throughout. The impact of nonzero vibrational linewidths on the presented results is described in Results and Discussion.
For the sake of simplicity, we assumed the thermalization of the background radiation is faster than all considered radiative transitions. The rate of change of the population in the $i$th vibrational state ($N_{i}$) is given by
@@
where $k_{ij}$ denotes rate of absorption when $i>j$ and emission when $i<j$ and $i_{\text{max}}$ is the highest energy bound state (Fig. 1a). This state is only loosely bound and is assumed to undergo irreversible dissociation at a rate $k_{\text{loss}}\gg k_{ij}\,\forall~{}i,j$. Equation 1 can be written in matrix notation as
@@
where N is a column vector containing the vibrational populations and $\mathbf{J}$ is the transition matrix. The lowest eigenvalue $\lambda_{0}$ of J is the rate constant of the unimolecular reaction [32, 33] of the described mechanism. Note that this mechanism does not consider dissociation from intermediate vibrational levels ($\nu<i_{\text{max}}$). Thus, our results are expected to be valid for BIRD promoted by multiphoton dissociation through a particular doorway state, which is the highest-energy bound state in our case.
Radiative transition rates
Assuming the molecular system that undergoes BIRD is weakly coupled to the radiation field, the transition rates $k_{ij}$ can be obtained from perturbation theory (Fermi’s Golden Rule [34, 35]) and written in terms of Einstein coefficients in free space [36, 37] and their generalizations in a weakly coupled and polaritonic microcavity. Rotational dynamics is much faster than thermal infrared radiative transition, so we employ isotropically averaged radiative rates in the long-wavelength limit (electrical dipole approximation) throughout this work.
@@
The final pieces necessary to compute the rate constants above are the transition dipole matrix $\mu_{ij}$ and the frequencies $\omega_{ij}$. We estimate the latter assuming the electronic ground-state potential energy curve of the examined diatomic molecules is given by the Morse potential [Fig. 1(a)][39]. The bound wave functions associated with the Morse potential are also employed in the computation of the vibrational transition dipole moments $\mu_{ij}$ via
@@
We investigated the IR radiative dissociation of hydrogen fluoride, HF, and lithium hydride, LiH. The employed dipole function $\mu(r)$ for HF is $\mu(r)=\mu_{0}re^{-\zeta r^{4}}$, where $\mu_{0}=0.45414e$ and $\zeta=0.081616$-4 were obtained from previous work on laser-induced HF dissociation[41]. The LiH dipole function was interpolated from ab initio quantum chemical computations (CCSD(T)/aug-cc-pVTZ, see Supplementary Information for further details). Morse potential parameters include the equilibrium bond length ($r_{e}$), electronic ground-state dissociation energy ($D_{e}$), and harmonic frequency ($\omega_{e}$) of each molecule and are listed in Table 1.
Table 1: Parameters used for Morse potential of HF [42, 43, 44, 41] and LiH [45].
Constant HF LiH
$D_{e}$ (eV) 6.123 2.641
$r_{e}$ (Å) 0.926 1.595
$\nu_{e}$ (cm-1) 4138 1405
$\nu_{e}\chi_{e}$ (cm-1) 86.7 28.3
$m_{r}$ (a.u) 0.95 0.88
The examined dissociation processes (via the most energetic bound state) have reduced room temperature rates relative to typical BIRD experiments involving polyatomic ions [27, 46, 29, 22] due to the relatively high stability and restricted number of state-space pathways available for diatomic dissociation proceeding via the highest energy level. This feature motivated our choice of examining diatomic BIRD at temperatures $T=\mathcal{O}(\hbar\omega_{e}/k_{\text{B}})$, an order of magnitude greater than room temperature. The SI shows that the trends discussed below hold for a wide interval of temperatures. Despite their different energetics, microcavity effects on HF and LiH BIRD yield the same qualitative behavior.
The transition rates $k_{ij}$ are proportional to the electromagnetic density of states $D(\omega)$, suggesting BIRD rates can be controlled by placing the reactive molecule inside a microcavity. To examine this scenario, we adapt the free space formalism described above to arbitrary photonic devices by replacing $D_{0}(\omega)$ in Eqs. 3 and 5 with the appropriate DOS for a weakly coupled microcavity $D_{\text{C}}(\omega)$.
@@
where $\lfloor x\rfloor$ denotes the floor function.
Lastly, we considered the BIRD scenario in which the diatomic interacts weakly with the electromagnetic component of polariton modes emergent from vibrational strong coupling between the microcavity modes and a material system (solid-state or molecular ensemble) with sufficiently large IR oscillator strength. We emphasize this material system does not undergo BIRD; its role is to generate polariton normal-modes with a tunable density of states and thermal energy density, as shown in Figs. 1(d) and (e). The formalism discussed above for BIRD in a weakly coupled microcavity can be straightforwardly adapted to model polariton-mediated BIRD (see SI) by replacing the photon density of states $D_{0}(\omega)$ with the photon-weighted polariton density of states $D_{P}(\omega)$.
@@
$D_{P}(\omega)=\sum_{m=0}^{\infty}\left(1-\frac{\delta_{0,m}}{2}\right)\frac{q(m,\omega)P_{C}(\omega)}{\pi v_{g}(\omega,q)L_{C}}\theta[\omega_{\text{C}}(\omega)-\frac{m\pi c}{L_{C}}],$ (11)
where $\theta(x)=1$ for $x\geq 0$ and vanishes elsewhere, $\omega_{\text{C}}(\omega)$ is the frequency of the photon mode that under strong coupling leads to the formation of a polariton with frequency $\omega$, $P_{C}(\omega)$ is the frequency-dependent photon content of polaritons with frequency $\omega$
@@
Our model shows singular behavior for $D_{P}(\omega)$ as $\omega\rightarrow\omega_{M}$. In this limit, both the photon content $P_{C}(\omega)$ and the group velocity $v_{g}(\omega,q)$ approach zero, but the ratio $P_{C}(\omega)/v_{g}(\omega,q)\rightarrow\infty$ resulting in $D_{P}(\omega)$ diverging to positive infinity. This singularity is due to the formation of the well-known stopgap region[59, 60, 61, 62] where $D_{P}(\omega)=0$. As previous work shows [63], introducing an ultraviolet cutoff and material disorder eliminates the singularity of $D_{P}(\omega)$ at $\omega_{M}$. Nevertheless, there remains significant enhancement of $D_{P}(\omega)$ in a neighborhood of $\omega_{M}$, which depends in a non-universal way on the type of disorder and coherence loss mechanism. This motivates our decision to use the formally exact expression presented in Eq. 11 and interpret our results as upper bounds on the effect of polaritons on BIRD.

### v3

@@ Abstract
Vibrational strong light-matter coupling offers a promising approach for controlling chemical reactivity with infrared microcavities. While recent research has examined potential mechanisms for this phenomenon, many important questions remain, including what type of reactions can be modified and to what extent this modification can be achieved. In this study, we explore the dynamics of Blackbody Infrared Radiative Dissociation (BIRD) in microcavities under weak and strong light-matter interaction regimes. Using a Master equation approach, we simulate the effects of infrared field confinement and polariton formation on BIRD rates for diatomic molecules weakly coupled to the radiation field. We present a framework explaining how infrared microcavities influence BIRD kinetics, highlighting the importance of overtone transitions in the process. Our findings outline conditions under which significant enhancement or mild suppression of BIRD rates can be achieved, offering insights into practical limitations and new strategies for controlling chemistry within infrared resonators.
@@ I   Introduction
Chemical reaction control is critical in applications ranging from industrial catalysis to medicine. Conventional methods for manipulating reaction rates involve varying temperature, pressure, or catalyst composition. An unconventional approach for controlling thermal reaction kinetics has been recently reported [1, 2, 3, 4] employing infrared (IR) microcavities [5] consisting of two moderate-quality mirrors separated by a distance $L_{C}$ of the order of IR wavelengths resonant with typical molecular vibrations [6]. When a collection of molecules with sufficiently large IR oscillator strength is embedded in a resonant microcavity, vibrational strong coupling occurs as signaled by the emergence of hybrid quasiparticles (vibrational polaritons) consisting of a superposition of molecular and electromagnetic modes of the confined device [7, 8, 9]. Recent theoretical research has probed potential mechanisms for the observed polariton effects on chemical reaction rates[10, 11, 12, 13, 14, 15, 16, 4, 17, 18, 19], and experimental progress towards the characterization of gas-phase reactivity under strong light-matter coupling has also been reported [20, 21]. Still, several fundamental questions remain[22, 23], including what types of reactions can be controlled with an optical microcavity and to what extent this control can be exerted. In this study, we investigate gas-phase unimolecular dissociation activated by the absorption of microcavity blackbody infrared radiation [24] under weak and strong coupling with a suitable material.
@@
Blackbody infrared radiative dissociation (BIRD) was proposed by Perrin in 1913 as a mechanism for gas phase unimolecular dissociation [25]. However, experimental observations in the 1920s ruled BIRD out in favor of the collisional activation mechanism [26, 27, 28]. Several decades later, the BIRD hypothesis was proven appropriate at sufficiently low pressures[29, 30, 31]. Since then, BIRD has become a valuable tool for investigating the thermal dissociation kinetics of molecular ions and clusters and has been observed in several systems, including ion clusters, transition-metal complexes, and biopolymers [24].
@@
The thermal radiation spectrum heavily influences BIRD rates. Given that free space and microcavity electromagnetic density of states and thermal radiation density can be significantly different [5], microcavity BIRD is expected to proceed with different rates compared to free space BIRD. However, the extent of this change, whether suppression or enhancement can be achieved, and if strong light-matter coupling introduces new possibilities relative to the weak coupling regime remain open questions. In this study, we address these questions by examining the dissociation of diatomic molecules driven by multiphoton or multipolariton absorption in microcavities. The qualitative conclusions drawn from this analysis are expected to be generalizable to polyatomic systems.
The primary aim of this work is to characterize thermal radiative dissociation in microcavities. Whereas prior studies employed hierarchical equations of motion and analytical rate theories based on Fermi’s golden rule to explore energy-diffusion-limited reactivity modeled by double-well potentials [14, 16, 32, 33, 34, 17], here we employ a Pauli master equation to investigate irreversible bond breaking with a Morse potential in a scenario where the microcavity or polaritons act solely as passive modifiers of the electromagnetic environment. Our methodology further differs by its microscopically detailed treatment of mechanical and electrical anharmonicity inherent to reactivity processes, and by its inclusion of the coupling of both fundamental and overtone vibrational transitions to the modified photon density of states, thereby providing a detailed description of the microcavity effect on the multiple reactive pathways available in BIRD.
@@
We employ a Pauli Master equation to investigate BIRD in (i) free-space, (ii) under weak coupling with a nearly empty microcavity, and (iii) under weak coupling with a polaritonic material. Our results give upper bounds for BIRD enhancement and suppression in weakly and strongly coupled microcavities with perfect mirrors (see SI Sec. 8 for a discussion of leaky microcavities) and identify overtone transitions as crucial ingredients for the observed modulation of BIRD. In Sec. II, we describe our methodology, Sec. III provides the main results and related discussion, and Sec. IV summarizes our main results and conclusions.
@@ II   Methods
Figure 1: (a) Illustration of the Morse potential describing the vibrational states and considered transitions of the NaLi molecule. (b) Illustration of a diatomic molecule inside a high quality planar microcavity with longitudinal length $L_{C}$. (c) Dispersion curves of transverse-electric (TE) and transverse-magnetic (TM) microcavity modes. (d) Scheme illustrating a reactive diatomic molecule inside a planar resonator strongly coupled to a host material system with significant oscillator strength and transition frequency $\omega_{M}$ (dashed line at $\omega=0.05~{}\text{eV}$). In this setup, the diatomic dissociation is mediated by absorption of thermal polaritons. (e) Illustration of polariton dispersion curves.
II.1      Pauli Master Equation for diatomic BIRD
Our analysis of BIRD in microcavities employs the same kinetic framework that has proven successful for BIRD studies in free space [30, 35, 36, 24]. Specifically, we propagate the populations of bound vibrational levels with a Pauli master equation[37] including transition rates obtained from Fermi’s golden rule [38]. This choice is justified because the infrared absorption and emission events of interest ($>$ 100 ns) are orders of magnitude slower than rovibrational decoherence, which randomizes vibrational coherences within on a much faster timescale. The resulting separation of time scales renders the diatomic reduced density matrix effectively diagonal (in the Morse potential Hamiltonian eigenstate basis) well before any population transfer occurs, allowing the Liouville–von Neumann equation to be coarse-grained to the Pauli form [37, 39, 40].
We neglect collisional activation and assume that dissociation is driven exclusively by photon absorption, consistent with the conditions of BIRD experiments[30, 24]. In such experiments, the use of dilute gas samples at low pressures ensures that intermolecular collisions are negligible. External fields are applied to trap and stabilize molecular trajectories within the interior of the reaction vessel. This prevents molecules from exchanging energy with walls and allows radiative processes to dominate [41, 42, 43]. Rotational degrees of freedom are ignored since they are weakly coupled to vibrational transitions, and the rotational dynamics are essentially classical at the temperatures of interest to BIRD. We also neglect Doppler and lifetime broadening throughout.
For the sake of simplicity, we assumed that the thermalization of the background radiation is faster than all the considered radiative transitions. We analyzed microcavity effects on BIRD occurring via two distinct mechanisms. In the first, denoted, dissociation via the doorway state, the dissociation proceeds only via the highest-energy bound state (the doorway state). In this case, the rate of change of the population in the $i$th vibrational state ($N_{i}$) is given by
@@ II.1      Pauli Master Equation for diatomic BIRD
where $k_{ij}$ denotes rate of absorption when $i>j$ and emission when $i<j$ and $i_{\text{max}}$ is the highest energy bound state (Fig. 1a). As mentioned, this state is only loosely bound and is assumed to undergo irreversible dissociation at a rate $k_{\text{loss}}\gg k_{ij}\,\forall~{}i,j$. Equation 1 can be written in matrix notation as
@@
where N is a column vector containing the vibrational populations and $\mathbf{J}$ is the transition matrix. The lowest eigenvalue $\lambda_{0}$ of J is the rate constant of the unimolecular reaction [44, 45] of the described mechanism. Note that this mechanism does not consider dissociation from intermediate vibrational levels ($i<i_{\text{max}}$). Thus, our results are expected to be valid for BIRD promoted by multiphoton dissociation through a particular doorway state, which in our case is the highest-energy bound state.
The second dissociation mechanism considered allows for direct radiative dissociation from any bound state of the diatomic molecule. Compared to Eq. 1, this introduces a dissociation channel for each bound state, characterized by a loss rate $k^{i}_{\text{loss}}$. The corresponding population rate equations are given by
$\frac{dN_{i}(t)}{dt}=\sum_{j\neq i}^{\nu_{\text{max}}}\left[k_{ij}N_{j}(t)-k_{ji}N_{i}(t)\right]-k^{i}_{\text{loss}}N_{i}(t).$ (3)
The transition matrix $\mathbf{J}$ now includes contributions from bound-continuum decay rates in all diagonal elements. The presence of $k_{i}^{\text{loss}}$ for all $i$ ensures that the vibrational population can be irreversibly removed from each level due to dissociation. The matrix $\mathbf{J}$ remains positive-definite, and its smallest eigenvalue $\lambda_{0}$ continues to govern the long-time dynamics of the system.
II.2      Radiative transition rates
Assuming the molecular system that undergoes BIRD is weakly coupled to the radiation field, the transition rates $k_{ij}$ can be obtained from perturbation theory (Fermi’s Golden Rule [46, 47]) and written in terms of Einstein coefficients in free space [48, 49] and their generalizations in a weakly coupled and polaritonic microcavity. Rotational dynamics is much faster than the thermal infrared radiative transition, so we employ isotropically averaged radiative rates in the long-wavelength limit (electrical dipole approximation) throughout this work.
@@ II.2      Radiative transition rates
The final pieces necessary to compute the rate constants above are the transition dipole matrix $\mu_{ij}$ and the frequencies $\omega_{ij}$. We estimate the latter assuming the electronic ground-state potential energy curve of the examined diatomic is given by the Morse potential [Fig. 1(a)][51]. The bound wave functions associated with the Morse potential are employed in the computation of the vibrational transition dipole moments $\mu_{ij}$ via
@@
The rates $k_{\text{loss}}^{i}$ for radiative absorption into the continuum by each bound state $i$ are obtained by a simple generalization of Eqs. 5 as discussed in the Supplementary Material.
We investigated the thermal IR radiative dissociation of the NaLi molecule. The electrical dipole function and Morse potential were obtained via interpolation from previous CCSDT/cc-pCVQZ calculations available in the literature [53] (see Supplementary Information for further details). Morse potential parameters are listed in Table 1.
Table 1: Parameters characterizing the Morse potential associated to NaLi.
Parameter Notation (unit) Value
Electronic ground-state dissociation energy $D_{e}$ (eV) 0.882
Equilibrium bond length $r_{e}$ (Å) 2.895
Harmonic frequency $\omega_{e}~{}\left(\text{cm}^{-1}\right)$ 257.4
Anharmonicity $\omega_{e}\chi_{e}~{}\left(\text{cm}^{-1}\right)$ 2.33
Reduced mass $m_{r}$ (a.u) 5.331
The transition rates $k_{ij}$ are proportional to the electromagnetic density of states $D(\omega)$, suggesting BIRD rates can be controlled by placing the reactive molecule inside a microcavity. To examine this scenario, we adapt the free space formalism described above to arbitrary photonic devices by replacing $D_{0}(\omega)$ in Eqs. 4 and 6 with the appropriate DOS for a weakly coupled microcavity $D_{\text{C}}(\omega)$.
@@
where $\lfloor x\rfloor$ denotes the floor function. This work focuses on establishing an upper limit for microcavity effects on BIRD by assuming perfectly reflecting mirrors. Real microcavities are leaky, so to assess the robustness of our predictions, we also computed the photon density of states and corresponding quantum state transition rates in imperfect microcavities with finite reflectivity and absorbing metallic mirrors using the electromagnetic field dyadic Green function [56] (see Supplementary Information for details).
Lastly, we considered the BIRD scenario in which the diatomic interacts weakly with the electromagnetic component of polariton modes emergent from vibrational strong coupling between the microcavity modes and a material system (solid-state or molecular ensemble) with sufficiently large IR oscillator strength. We emphasize that the polariton modes arise from collective strong coupling between the cavity and a host material with substantial infrared oscillator strength. The reactive molecule experiences modified radiative dynamics due to changes in the photon density of states induced by the formation of polaritons, as illustrated in Figs. 1(d) and (e). Consequently, in this study, the interaction strength between our reactive diatomic and the radiation field is obtained from first principles without any scaling parameters.
While this polariton-assisted scenario poses challenges for experimental realization as BIRD generally requires low-density, collision-free conditions, it is introduced here as a theoretical construct that isolates the effect of polariton-modified photonic environments on molecular infrared radiative dissociation. Nonetheless, one could envision suppressing collisions by confining reactive molecules using transparent partitions or external fields, while maintaining exposure to the confined electromagnetic modes[41, 42, 43]. A detailed treatment of these experimental considerations, as well as the role of direct interactions between host and reactive molecules, lies beyond the scope of the present work and warrants future investigation.
The formalism discussed above for BIRD in a weakly coupled microcavity can be straightforwardly adapted to model polariton-mediated BIRD (see SI) by replacing the photon density of states $D_{0}(\omega)$ with the photon-weighted polariton density of states $D_{P}(\omega)$.
@@
$D_{P}(\omega)=\sum_{m=0}^{\infty}\left(1-\frac{\delta_{0,m}}{2}\right)\frac{q(m,\omega)P_{C}(\omega)}{\pi v_{g}(\omega,q)L_{C}}\theta[\omega_{\text{C}}(\omega)-m\pi c/L_{C}],$ (12)
where $\theta(x)=1$ for $x\geq 0$ and vanishes elsewhere, $\omega_{\text{C}}(\omega)$ is the frequency of the photon mode that under strong coupling leads to the formation of a polariton with frequency $\omega$, $P_{C}(\omega)$ is the photon content of the polariton modes with frequency $\omega$
@@
Our model shows singular behavior for $D_{P}(\omega)$ as $\omega\rightarrow\omega_{M}$. In this limit, both the photon content $P_{C}(\omega)$ and the group velocity $v_{g}(\omega,q)$ approach zero, but the ratio $P_{C}(\omega)/v_{g}(\omega,q)\rightarrow\infty$ resulting in $D_{P}(\omega)$ diverging to positive infinity. This singularity is due to the formation of the well-known stopgap region[66, 67, 68, 69] where $D_{P}(\omega)=0$. As previous work shows [70], introducing an ultraviolet cutoff and material disorder eliminates the singularity of $D_{P}(\omega)$ at $\omega_{M}$. Nevertheless, there remains significant enhancement of $D_{P}(\omega)$ in a neighborhood of $\omega_{M}$, which depends in a non-universal way on disorder and coherence loss mechanisms. Similarly, the stopgap is known to contain weakly coupled states in disordered systems causing $D_{P}(\omega)$ to become nonzero in this region. However, the photon density of states remains lower in the stopgap in comparison to the weak coupling regime.
The robustness of the enhanced and reduced photon-weighted polariton density of states at frequencies $\omega$ approaching $\omega_{M}$ from the left and right, respectively, support our choice to to use the formally exact expression presented in Eq. 12 and interpret our results as upper bounds on the effect of polaritons on BIRD. The relevance of these bounds to experiments will be shown to depend on the collective light-matter interaction strength and the detuning between transitions of interest and $\omega_{M}$ as discussed in detail in Results and Discussion.